OpenStudy (anonymous):

what is sin 2pi/3? how do you solve wihtout a calculator?

OpenStudy (anonymous):

Root3/2.

OpenStudy (anonymous):

but how do i solve it? i completely forgot trig

myininaya (myininaya):

unit circle

OpenStudy (anonymous):

2pi/3 is the same at pi/3. pi/3 is a 60 degree angle. Sin is Opposite/Hypotenuse. so it would be Root3/2.

myininaya (myininaya):

\[\frac{2 \pi}{3} \cdot \frac{180^o}{\pi}=120^o\] maybe you think better using degrees